# View Selection for 3D Captioning via Diffusion Ranking

### v1

@@ 4 Dataset
In this section, we detail our process for correcting the Cap3D captions, expanding the dataset with high-quality 3D objects from Objaverse-XL, and ethical filtering. More detailed hyper-parameters and comparisons are included in Appendix B.
@@ 4.1 Correction of Cap3D Captions
As Cap3D contains a lot of good quality captions as shown in their paper and public dataset, our first objective is to identify erroneous Cap3D captions, which might contain incorrect information or hallucinations. We tried three strategies as outlines the below.
Image-Text Alignment Method: We discovered that utilizing the maximum and average CLIP scores effectively filters out inaccurate captions. Most of erroneous captions, like those depicted in Figure 1, described improbable combinations of objects (e.g., “a mix of a frog, teddy bear, and monster” or “an orangutan accompanied by a pelican and a fish”) in scenarios where only one entity was present in the given 3D object. Such discrepancies arise when different views of the same 3D object receive varied entity captions from BLIP2, which GPT4 then erroneously combines, shown in Figure 3. To detect this kind of case, we computed both the average and maximum CLIP scores between the final caption and all eight rendered views used in Cap3D. A validation set of objects with inaccurate captions was annotated and used to determine two thresholds (mean & max as shown in Figure 5), with the goal of encompassing all objects in this set. We then use the two selected thresholds to filter out possible issued objects out of a total of .
@@
Text-Based Method: Attempting to identify errors solely based on captions proved challenging due to the diverse and complex nature of objects within Objaverse, making it difficult to detect hallucinations based on text alone. This complexity arises because some 3D objects genuinely comprise multiple or unusual components. Despite this, we developed a technique for identifying the misuse of terms related to “image” and “rendering”, as these are directly associated with the rendering process rather than the 3D objects themselves. Through this method, we identified approximately 23,000 objects requiring correction.
@@ 4.2 Dataset Expansion and Ethical Filtering
Our expansion includes adopting the remaining objects of Objaverse, where Cap3D did not include, and high-quality 3D objects from Objaverse-XL’s curated subset (Section 4.1 of Deitke et al. (2023b), selected through human evaluation and heuristics from a pool of 10 million objects. This extension enhances the diversity and quality of our dataset. For the detailed object uids, please refer to the CSV file we attached in appendix.
@@
We compared caption length and n-grams Bird et al. (2009) of captions among Human, Cap3D, and our captions in a 5k common set. As shown in Figure 7, our captions usually contain longer length indicating more details than Cap3D and human-authored captions. Table 7 demonstrates we have the largest vocabulary size.
@@ 5 Experiments
In this section, we compare our captions against Cap3D captions and human-authored captions in terms of quality and hallucination degrees through human studies. We also ablate our methods to verify the effectiveness of the proposed DiffuRank. Then, we compare text-to-3D models finetuned on Cap3D and our updated Captions on the same set to measure the improvements of caption alignment at scale. Finally, we further verify the effectiveness of our propose DiffuRank by examining it on a VQA task. For the sake of space, we list quantitative results here and include qualitative comparisons in Appendix B.3, B.4, B.5, and C.
@@
5.1 Captioning Evaluation
Metrics. Our primary evaluation method utilizes A/B testing with human judgment, where participants evaluate a pair of captions on a 1-5 scale, with 3 representing a neutral preference (i.e., tie). Our approach includes two distinct assessments: (a) evaluating which caption more accurately describes the object’s type, appearance, and structure, and (b) determining which caption is less prone to presenting incorrect information or hallucinations. Each assessment involves over 10,000 ratings across 4,000 objects to ensure statistical reliability. We calculate and report the average scores and the frequency each option is preferred (i.e., excluding neutral (tie) responses). More human evaluation details are included in Appendix B.7. Additionally, we follow Cap3D Luo et al. (2023a) and employ automated metrics, including CLIP score, measuring the cosine similarity between CLIP encodings and input images, and CLIP R percision Poole et al. (2022), assessing the match between a rendered image and all potential texts.
@@ 5.1 Captioning Evaluation
A comparison of our method, which selects the top-6 views, with alternatives—the bottom-6 views and horizontally placed 6-views—demonstrates the impact of DiffuRank on performance. Specifically, as depicted in Figure 2, bottom-6 views often relate less to the 3D object as they may capture only the back or bottom. This issue highlights the difficulties arising from Objaverse’s random default orientation, positioning cameras ‘horizontally’ does not always ensure they are actually horizontal. More qualitative compairsons between the three types of view selection are included in Appendix B.5. Furthermore, DiffuRank does not consistently achieve optimal performance, as illustrated by the selection of the 6th image in the first row (referenced in Figure 2) captioned ’a blue laptop.’ Enhancements could be achieved through using an improved text-to-3D diffusion models, a topic explored in detail in Section 6.
@@ 5.2 Text-to-3D Generation with New Captions
Settings. This section we finetune Text-to-3D models to check if our updated captions can bring more improvements compared to Cap3D captions. For this purpose, we would mainly conduct experiments over point-E Nichol et al. (2022) and shap-E Nichol and Jun (2023) as they are used in Cap3D. We follow the same setting as Cap3D, including learning rate, batch size, optimizer, and steps. We adopted the same 330k training split and test split used in Luo et al. (2023a), and we have updated captions in this 330k set (20%). Additionally, we scale our experiment up, and train models with 825k ( data from our full 3D-text pairs. More details and qualitative results are included in Appendix C.
Metrics. We incorporated the use of CLIP Score and CLIP R-Precision Poole et al. (2022); Luo et al. (2023a) in our evaluation process. CLIP R-Precision involves ranking a rendered image among all text pairs within the test set based on the cosine similarity as measured by CLIP, then determining the precision based on accurate text-image matches. Given the availability of ground truth images, we employed the FID metric to compare the fidelity of 3D rendered images with these true images. Additionally, the evaluation included calculating the CLIP Score for these reference images.
Results. Results are showcased in Table 2. Considering we updated nearly 20% captions of the 330k training set for Cap3D 3D-text pairs, we anticipated some improvement, albeit modest. However, the improvements exceeded our expectations. Our enhanced model (‘model + Ours’ with 330K data points) consistently outperformed both the ‘model + Cap3D’ (with 330K data points) version and OpenAI’s pre-trained Shap·E models. Surpassing the OpenAI Shap·E models is non-trivial, as the ’model + cap3d’ version generally showed declining performance when compared to the pre-trained model. The performance enhancement achieved by correcting 20% of the data underscores the effectiveness of addressing misalignments in the 3D-text of Cap3D by locating the potential errors and refining with our new captioning approach. Furthermore, by expanding our dataset by 2.5 times, we’ve boosted performance across multiple metrics and models. Given that OpenAI’s Shap·E model was trained on proprietary data, our findings suggest that our 3D-text dataset could be a competitive open-source alternative.
@@ 5.3 DiffuRank on VQA
Settings. We extend our DiffuRank to solve Visual Question Answering task, with the help of a pre-trained text-to-2D diffusion model Rombach et al. (2022). We list our detailed settings and the updated algorithm in Appendix D. We mainly compare to CLIP Radford et al. (2021) in terms of zero-shot VQA performance and test on the Multimodal Visual Patterns (MMVP) benchmark Tong et al. (2024), comprising nine fundamental visual patterns across 150 images pairs. Each pair of images (Figure 8), despite having clear visual distinctions, are perceived similarly by the CLIP model. Each pair is associated with a question that has two divergent answers. Numerous Vision-Language Models (VLMs) have been shown to underperform on this challenging benchmark.
Given that the task involves Visual Question Answering (VQA), neither our approach nor the CLIP model is inherently designed to generate textual responses directly. To address this, we employed GPT-4 to transform each question and its corresponding answers into declarative statements. Consequently, for each image pair, we obtained two distinct statements. For DiffuRank, we executed multiple iterations of alignment estimation for the statements corresponding to each image, selecting the statement with the highest alignment estimate as the correct answer/statement. For CLIP model, we determined the appropriate answer by calculating the cosine similarity between an image and each statement, choosing the statement with the greatest similarity as the response. We used “ViT-B/32” CLIP here for evaluation.
Metrics. Our evaluation metrics are aligned with those proposed by Tong et al. (2024). A model’s response is deemed accurate only if it correctly identifies the appropriate statements for both images in a pair. Hence, if a model accurately selects the correct statement for only one image within the pair, its attempt is marked as incorrect. It is important to note that both DiffuRank and CLIP may occasionally select identical statements for different images within the same pair.
@@
6 Future Work & Limitations
Future Work: DiffuRank leverages a pre-trained text-to-3D diffusion model for rendering view ranking, enhancing 3D object captioning. Improved captioning enables the refinement of the diffusion model, creating a feedback loop that cyclically utilizes the model for data generation and employs this data to fortify the model further. Besides, due to our limited computational resources and funding, it is not feasible to encompass all Objaverse-XL objects, presenting an opportunity for industrial entities.
Limitations: During our subtitling process, we use DiffuRank to select 6 rendered views out of 28 views. This process requires us to render more views, generate captions, and perform inference using a pre-trained text-to-3D diffusion model to compute alignment scores. All of the steps take calculation and time.
As highlighted in the related work (Section 2), DiffuRank faces challenges with speed, requiring multiple samplings for each option and necessitating forward model processing for all options. Our process for a single 3D object involves 28 rendered views, 5 captions per view, and performing sampling 5 times ( in Alg. 1), resulting in a total of 700 inference operations. While parallel processing (large batch size) can mitigate delays, the system’s performance is inherently slow. We show a VQA extension in Section 5.3 as it only has two options. But, generally, DiffuRank’s design is not optimal for tasks requiring numerous options, such as classification and image-text retrieval.
@@ 8 Acknowledgement
This work has been made possible through the generous support of the “Efficient and Scalable Text-to-3D Generation” grant from LG AI Research, and the National Science Foundation (NSF) under Grant No. 1453651. We greatly appreciate Chris Rockwell for his invaluable technical support in caption evaluation, and Mohamed El Banani for his insightful feedback to our initial draft. Tiange thanks Minghua Liu and Jiaming Song for their insightful discussions back at NeurIPS 2023 in NOLA.
A Broader Impact
B Dataset: more details & results
B.1 Extra dataset details
B.2 Captions: overcome the failure cases in Cap3D
B.3 Captions: Ours vs. Cap3D vs. human-authored
B.4 Captions: Ours vs. ablated variants
B.5 Diffu: Ours vs. bottom 6-views vs. horizontal 6-views
B.6 Failure cases
B.7 Human evaluation details
C Text-to-3D: more details & results
C.1 Setting
C.2 Qualitative comparisons
D DiffuRank on VQA
Appendix A Broader Impact
By enhancing the accuracy and richness of captions for 3D objects, this work facilitates advancements in 3D modeling and prompote related applications in educational tools, interactive learning environments, and assistive technologies, making digital content more accessible and informative. Moreover, by addressing inaccuracies and hallucinations in captions which could be used in AI-content generations, our work underscores the pursuit of more reliable and trustworthy AI systems. During the process, we undertaken with a commitment to ethical considerations to filter out potential ethical issued 3D objects. We recognize the wide-reaching effects of our work on society and maintain that it chiefly offers positive contributions towards the progress of generative modeling and its implementation in diverse fields.
Appendix B Dataset: more details & results
B.1 Extra dataset details
In Section 4, we addressed approximately caption corrections for the Cap3D dataset, significantly reducing its hallucinations. Our efforts also expand the dataset to include over 1 million 3D-text pairs, encapsulating the entirety of the Objaverse Deitke et al. [2023a] and portions of the Objaverse-XL high-quality set Deitke et al. [2023b]. The objects with updated captions are cataloged in a CSV file within the supplementary material, accessible via “uid” or cryptographic hash values (”sha256”). These identifiers correspond to the ones provided in the Objaverse and Objaverse-XL datasets.
As mentioned in the Introduction, we are excited to also provide access to rendered images associated with each object. These images include detailed camera information (both intrinsic fov and extrinsic RT matrix), depth map, and MatAlpha, in addition to point clouds that complement the textual captions. Alongside these resources, we are releasing the source code for our DiffuRank methodology, which facilitates the replication of our findings. The distribution also includes pre-trained models, further aiding in the exploration and utilization of our dataset. This comprehensive package aims to empower researchers in our community. They will be released under ODC-By 1.0 license.
Our GPT4-Vision prompt is defined as “Renderings show different angles of the same set of 3D objects. Concisely describe 3D object (distinct features, objects, structures, material, color, etc) as a caption” accompanied by six image tokens. On average, the context encompasses approximately 1,867 tokens, while the average number of tokens generated stands at approximately 26.72. Notably, we employed the ”GPT-4-1106-vision-preview” model for this study.
As described in Section 3.3, given a 3D object, we generate 28 views using two distinct rendering methods Luo et al. [2023a], Jun and Nichol [2023]. For each view, we generate 5 captions with BLIP2. Subsequently, we apply the DiffuRank algorithm (Algorithm 1) to evaluate the alignment of the 28 renderings relative to the input 3D object by doing inference ovew 140 captions and the 3D object. Ultimately, we select the best 6 views for further caption generation using GPT4-Vision.
For the ray-tracking render engine, we used Blender render engine ‘CYCLES’ with samples 16. Additionally, we adopted ‘OPTIX’ denoiser for the cycle engine. For the real-time render engine, we used Blender render engine ‘EEVEE’ with ‘taa_render_samples’ 1.
B.2 Captions: overcome the failure cases in Cap3D
The Cap3D captions we used to compare thoroughout the whole paper are from their dataset page. Specically, the version described in their paper.
Here, we provide direction comparisons with the failure cases mentioned in their paper “Limitations and Failure Cases”. Our captions have obviously eliminated lots of hallucinations, such as ‘butterfly’ and ‘flowers’ in Figure 9, and ‘dump truck’ in Figure 10.
B.3 Captions: Ours vs. Cap3D vs. human-authored
We present a variety of qualitative comparisons: those generated by our model, those produced by Cap3D, and captions written by humans, all of which were selected through random sampling. The below qualitative results show the captions generated by our method usually contain more details and less hallucinations.
B.4 Captions: Ours vs. ablated variants
We list several qualitative comparisons here to demonstrate the effectiveness of our method compared to (1) Bottom 6-views, we employ the 6 renderings identified as having the lowest alignment scores, as determined by our DiffuRank algorithm (refer to Alg. 1); (2) Allviews 28-views, which involves utilizing all 28 rendered views as inputs for the GPT4-Vision; and (3) Horizontal 6-views, this configuration involves selecting 6 rendered views that position the camera horizontally relative to the object’s default orientation, adhering to the same vertical positioning guidelines used by Cap3D. Results generally show the captions generated by our method (i.e., Top-6) contain more accurate, detailed, and less hallucinated information.
B.5 Diffu: Ours vs. bottom 6-views vs. horizontal 6-views
This section lists several randomly sampled DiffuRank results of Top 6-views with 6 highest alignment scores (our method), bottom 6-views, and horizontal 6-views. According to the results, we can see (1) Top 6-views obviously outperforms Bottom 6-views on Figures 57, 58, 59, 61, 66, 67, 68, 71; (2) Compared to Horizontal 6-views, DiffuRank can adaptly choose angles and types of rendering as shown in Figures 55, 70, 72; (3) in some cases (Figures 54, 63, 65), there are no significant difference.
B.6 Failure cases
We have observed three types of failure cases: (1) DiffuRank fails due to BLIP2 captioning fails or alignment compute not accurate. As shown in Figure 77, where BLIP2 captions contain a lot of “a tree in the dark”. Since our DiffuRank needs the initial captioning results to compute alignment scores, BLIP2 captioning fails will cause rendering selection poorly and further cause final caption inaccurate. This could be solved via stronger captioning model, such as GPT4-Vision. Also, as mentioned in Future work (Section 6), with better captions, we can fine-tune stronger Text-to-3D models, which help to obtain more accurate alignment scores. (2) sometimes, our captioning method fails to capture small object. One example is in Figure 9, where there is a small black person above the rock, while the caption fails to describe it. Also, it may contain hallucinations with small chances (according to our eyeballs over 10k captions) as shown in Figure 76. (3) for some scene renderings, the model failed to capture meaningful characteristics for Figure 78 with caption “Abstract 3D composition with fragmented, textured surfaces in shades of beige, white, and charcoal”. However, human may also not distinguish this kind of renderings.
B.7 Human evaluation details
We utilize the Hive platform for conducting crowdsourced A/B testing. In this process, participants are presented with an image accompanied by two different captions as shown in Figure 79. They are asked to judge which caption is more suitable based on a 5-point scale, where a score of 3 indicates neither caption is preferred over the other. Scores of 1 and 2 suggest a preference for the left caption, with 1 indicating a strong preference and 2 a moderate preference. The sequence in which the captions are presented (left or right) is varied randomly in each case.
Participants receive guidelines on how to perform this task, including examples that set the standard for quality. We have two distinct types of tasks as shown in Table 1: quality and hallucination. For quality tasks, workers are advised to focus first on the accuracy of their choices, followed by the level of detail provided in terms of type, structure, and appearance. For hallucination tasks, workers are advised to focus on if the caption contain hallucination or false information.
We totally hired 46 workers from Hive without access to their personally identifiable information. They are paid approximately $35 per 1k tasks for our caption evaluation tasks. The entire procedure was carried out in compliance with the ECCV ethics guidelines.
The platform automatically excludes workers who fail to meet the required standards on essential test examples set by us. However, our review revealed that some workers managed to meet the criteria for these essential examples but engaged in deceitful practices for the rest. The prevalent forms of deceit included consistently choosing the same option (always choose left or right) or selecting captions based on their length, either the shortest or the longest. Consequently, we conducted a thorough examination of all workers and excluded those found to be engaging in these deceptive practices, also disregarding their evaluations.
Appendix C Text-to-3D: more details & results
In this section, we provide a detailed examination of our Text-to-3D experiments, along with a comprehensive set of qualitative comparisons. It is important to note that employing captions generated by our method typically enhances the performance of Shap·E pre-trained models, a trend that is clearly supported by the data presented in Table 2. However, when we fine-tune the Shap·E pre-trained model using Cap3D, we observe a decline in performance across all CLIP-based metrics.
C.1 Setting
We adopted the same fine-tune strategy used in Cap3D Luo et al. [2023a] for fair comparisons. We employed the AdamW optimizer alongside the CosineAnnealingLR scheduler, setting the initial learning rate at for fine-tuning both the Point·E and Shap·E models. The batch sizes were set to for Shap·E and for Point·E. For training epochs, we set the training epoch which would cost approximately three days. The training was performed on four A40 GPUs.
The evaluation times, measured in seconds per iteration and inclusive of rendering, are as follows:
For Point·E, the total time is 37 seconds, with 28 seconds dedicated to text-to-3D processing and 9 seconds to rendering.
Shap·E (stf) requires 16 seconds in total for both text-to-3D processing and rendering.
Shap·E (NeRF) takes significantly longer, with a total of 193 seconds for both text-to-3D processing and rendering.
C.2 Qualitative comparisons
Appendix D DiffuRank on VQA
Algorithm 2 demonstrates the DiffuRank approach to the task of 2D Visual Question Answering. Initially, the process involves converting the question and each potential answer/option into a coherent statement. As shown in Figure 8, we convert Question: “Is the school bus driving towards or away from the camera?” and options “(a) Towards the camera (b) Away from the camera” into statements (1) “The school bus is driving towards the camera and statement” and (2) “The school bus is driving away from the camera”. Another example shows converting Question: “Is there a shadow on the flower?” and options “(a) Yes (b) No,(a)” into statements (1) “There is a shadow on the flower.” and (2) “There is not a shadow on the flower.”
This conversion is accomplished through the utilization of GPT-4 in our implementation. Subsequently, we determine the alignment scores by evaluating the correspondence between each generated statement and the provided 2D image. The statement that exhibits the highest alignment score, along with its associated option, is then selected as the definitive answer.
Different from Algorithm 1, our objective here is computed over noise difference, the way adopted in our used stable-diffusion models Rombach et al. [2022].

### v2

@@ 4 Dataset
This section details our process for correcting the Cap3D captions, expanding the dataset with high-quality 3D objects from Objaverse-XL, and ethical filtering. More detailed hyper-parameters and comparisons are included in Appendix B.
@@ 4.1 Correction of Cap3D Captions
As Cap3D contains a lot of good quality captions, our first objective is to identify erroneous Cap3D captions, which might contain incorrect information or hallucinations. We tried three strategies as outlines the below.
Image-Text Alignment Method: We discovered that utilizing the maximum and average CLIP scores effectively filters out inaccurate captions. Most of erroneous captions, like those depicted in Figure 1, described improbable combinations of objects (e.g., “a mix of a frog, teddy bear, and monster" or “an orangutan accompanied by a pelican and a fish") in scenarios where only one entity was present in the given 3D object. Such discrepancies arise when different views of the same 3D object receive varied entity captions from BLIP2, which GPT4 then erroneously combines, shown in Figure 3. To detect this kind of case, we computed both the average and maximum CLIP scores between the final caption and all eight rendered views used in Cap3D. A validation set of objects with inaccurate captions was annotated and used to determine two thresholds (mean & max as shown in Figure 5), with the goal of encompassing all objects in this set. We then use the two selected thresholds to filter out possible issued objects out of a total of .
@@
Text-Based Method: Attempting to identify errors solely based on captions proved challenging due to the diverse and complex nature of objects within Objaverse, making it difficult to detect hallucinations based on text alone. This complexity arises because some 3D objects genuinely comprise multiple or unusual components. Despite this, we developed a technique for identifying the misuse of terms related to “image" and “rendering", as these are directly associated with the rendering process rather than the 3D objects themselves. Through this method, we identified approximately 23,000 objects requiring correction.
@@ 4.2 Dataset Expansion and Ethical Filtering
Our expansion includes adopting the remaining objects of Objaverse, where Cap3D did not include, and high-quality 3D objects from Objaverse-XL’s curated subset (Section 4.1 of [12], selected through human evaluation and heuristics from a pool of 10 million objects. This extension enhances the diversity and quality of our dataset.
@@ 5 Experiments
In this section, we compare our captions against Cap3D captions and human-authored captions in terms of quality and hallucination degrees through human studies. We also ablate our methods to verify the effectiveness of the proposed DiffuRank. Then, we compare text-to-3D models finetuned on Cap3D and our updated Captions on the same set to measure the improvements of caption alignment at scale. Finally, we further verify the effectiveness of our propose DiffuRank by examining it on a VQA task. For the sake of space, we list quantitative results here and include qualitative comparisons in Appendix B and C.
5.1 Captioning Evaluation
Metrics. Our primary evaluation method utilizes A/B testing with human judgment, where participants evaluate a pair of captions on a 1-5 scale, with 3 representing a neutral preference (i.e., tie). Our approach includes two distinct assessments: (a) evaluating which caption more accurately describes the object’s type, appearance, and structure, and (b) determining which caption is less prone to presenting incorrect information or hallucinations. Each assessment involves over 10,000 ratings across 4,000 objects to ensure statistical reliability. We calculate and report the average scores and the frequency each option is preferred (i.e., excluding neutral (tie) responses). More human evaluation details are included in Appendix B.7. Additionally, we follow Cap3D [36] and employ automated metrics, including CLIP score, measuring the cosine similarity between CLIP encodings and input images, and CLIP R percision [46], assessing the match between a rendered image and all potential texts.
@@ 5.1 Captioning Evaluation
A comparison of our method, which selects the top-6 views, with alternatives—the bottom-6 views and horizontally placed 6-views—demonstrates the impact of DiffuRank on performance. Specifically, as depicted in Figure 2, bottom-6 views often relate less to the 3D object as they may capture only the back or bottom. This issue highlights the difficulties arising from Objaverse’s random default orientation, positioning cameras ‘horizontally’ does not always ensure they are actually horizontal. More qualitative compairsons between the three types of view selection are included in Appendix B.5. Furthermore, DiffuRank does not consistently achieve optimal performance, as illustrated by the selection of the 6th image in the first row (referenced in Figure 2) captioned ’a blue laptop.’ Enhancements could be achieved through using an improved text-to-3D diffusion models, a topic explored in detail in Appendix E.
@@ 5.2 Text-to-3D Generation with New Captions
Settings. This section we finetune Text-to-3D models to check if our updated captions can bring more improvements compared to Cap3D captions. For this purpose, we would mainly conduct experiments over point-E [42] and shap-E [41] as they are used in Cap3D. We follow the same setting as Cap3D, including learning rate, batch size, optimizer, and steps. We adopted the same 330k training split and test split used in [36], and we have updated captions in this 330k set (20%). Additionally, we scale our experiment up, and train models with 825k ( data from our full 3D-text pairs. More details and qualitative results are included in Appendix C.
Metrics. We incorporated the use of CLIP Score and CLIP R-Precision [46, 36] in our evaluation process. CLIP R-Precision involves ranking a rendered image among all text pairs within the test set based on the cosine similarity as measured by CLIP, then determining the precision based on accurate text-image matches. Given the availability of ground truth images, we employed the FID metric to compare the fidelity of 3D rendered images with these true images.
Results. Results are showcased in Table 2. Considering we updated nearly 20% captions of the 330k training set for Cap3D 3D-text pairs, we anticipated some improvement, albeit modest. However, the improvements exceeded our expectations. Our enhanced model (‘model + Ours’ with 330K data points) consistently outperformed both the ‘model + Cap3D’ (with 330K data points) version and pre-trained Shap·E model. Surpassing the pre-trained Shap·E model is non-trivial, as the ’model + Cap3D’ version generally showed declining performance when compared to the pre-trained Shap·E model, indicating that fine-tuning on Cap3D data actually harms the performance. The performance enhancement achieved by correcting 20% of the data underscores the effectiveness of addressing misalignments in the 3D-text of Cap3D by locating the potential errors and refining with our new captioning approach. Furthermore, by expanding our dataset by 2.5 times, we have boosted performance across multiple metrics and models. Given that Shap·E model was trained on proprietary data, our findings suggest that our proposed 3D-text dataset could be a competitive open-source alternative.
@@ 5.3 DiffuRank on VQA
Settings. We extend our DiffuRank to solve Visual Question Answering task, with the help of a pre-trained text-to-2D diffusion model [50]. We list our detailed settings and the updated algorithm in Appendix D. We mainly compare to CLIP [49] in terms of zero-shot VQA performance and test on the Multimodal Visual Patterns (MMVP) benchmark [60], comprising nine fundamental visual patterns across 150 images pairs. Each pair of images (Figure 8), despite having clear visual distinctions, are perceived similarly by the CLIP model. Each pair is associated with a question that has two divergent answers. Numerous Vision-Language Models (VLMs) have been shown to underperform on this challenging benchmark.
Given that the task involves Visual Question Answering (VQA), neither our approach nor the CLIP model is inherently designed to generate textual responses directly. To address this, we employed GPT-4 to transform each question and its corresponding answers into declarative statements. Consequently, for each image pair, we obtained two distinct statements. For DiffuRank, we executed multiple iterations of alignment estimation for the statements corresponding to each image, selecting the statement with the highest alignment estimate as the correct answer/statement. For CLIP model, we determined the appropriate answer by calculating the cosine similarity between an image and each statement, choosing the statement with the greatest similarity as the response. We used “ViT-B/32" CLIP here for evaluation.
Metrics. Our evaluation metrics are aligned with those proposed by [60]. A model’s response is deemed accurate only if it correctly identifies the appropriate statements for both images in a pair. Hence, if a model accurately selects the correct statement for only one image within the pair, its attempt is marked as incorrect. It is important to note that both DiffuRank and CLIP may occasionally select identical statements for different images within the same pair.
@@ Acknowledgement
This work has been made possible through the generous support of the “Efficient and Scalable Text-to-3D Generation" grant from LG AI Research. We also thank the OpenAI Researcher Access Program for partially supporting our use of the GPT-4 API. We greatly appreciate Chris Rockwell for his invaluable technical support in caption evaluation, and Mohamed El Banani for his insightful feedback to our initial draft. Tiange thanks Minghua Liu and Jiaming Song for their insightful discussions back at NeurIPS 2023 in NOLA.